# Condense to Conduct and Conduct to Condense

### v1

@@ 2.2 Proof of Theorem 2
We are to show that $\pi$ is a $(\alpha,\epsilon_{1},r)$-Multi-Source-Somewhere-Condenser. By definition, we should consider some independent random variables $U_{1},\ldots,U_{w}$ such that their min-entropy is at least $\alpha n$ and show that the distribution $\pi(U_{1},\ldots,U_{w})$ has the appropriate properties. However, without loss of generality (using Lemma 1), we can assume that the distribution of $U=U_{1}\times\ldots\times U_{w}$ is flat, that is, it is really just a uniformly distributed discrete $(2^{\alpha n})$-box $U$ (as defined in Definition 2). Then our original statement (after translating to the discrete interpretation) translates into the following statement:
@@
  • $|R|\leq r\cdot 2^{\alpha nw}$, where $r=w2^{-\alpha n\epsilon_{2}}+2^{-\alpha nw\epsilon_{3}}$;
@@
    $|C_{i}\cap\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{y\}}\times\ldots\times\{0,1\}^{n}|\leq 2^{-(1+\epsilon_{1})\alpha n}\cdot|C_{i}|.$
@@
We say that the set $X$ has at point $y\in\{0,1\}^{n}$, an $i$-Battleneck-Slice
@@
Intuitively: $i$-Battleneck-Slice at point $y$ means that the projection of $X$ onto the $i$-coordinate at point $y$ is nonempty, but suitably thin.
@@ 2. Step 2:
    If $\tilde{R}$ contains any (for any $i$ and at any point $y$) $i$-Battleneck-Slice $B$, perform the assignments:
@@ 3. Step 3:
    Loop Step 2 until $\tilde{R}$ no longer contains any Battleneck-Slice.
@@ 6. Step 6:
Intuitively: from the set $\pi(U)$ we cut off all Battleneck-Slices one by one until we are left with $R_{0}$, which no longer contains sets of this type. Then all cut off $i$-Battleneck-Slices concerning the $i$-th coordinate together form $C_{i}$, unless their measure is sufficiently small, in which case $C_{i}$ is empty and these Battleneck-Slices go to $R_{1}$. The final $R$ is the sum of $R_{0}$ and $R_{1}$.
@@
since every such set as above on the left is either empty or (as follows from the construction) equal to exactly one $i$-Battleneck-Slice cut off from $\tilde{R}$.
@@
We have to show $|R|\leq r\cdot 2^{\alpha nw}\text{, where }r=w2^{-\alpha n\epsilon_{2}}+2^{-\alpha nw\epsilon_{3}},$ that is,
$|R|\leq w2^{(w-\epsilon_{2})\alpha n}+2^{(1-\epsilon_{3})\alpha nw}.$
@@
$\begin{split}|R_{1}|&\leq w\cdot 2^{(w-\epsilon_{2})\alpha n},\\ |R_{0}|&\leq 2^{(1-\epsilon_{3})\alpha nw}.\end{split}$
@@
The last inequality follows directly from the fact that $R_{0}$ does not contain any Battleneck-Slice, the bound $|R_{0}|\leq|\pi(U)|=2^{\alpha nw}$ and the property (5) (this property holds for $\pi(U)$, so obviously also for $R_{0}\subset\pi(U)$), which, however, requires a slightly deeper analysis, which (thus concluding the entire proof) we present below:
@@
If the set $R_{0}$ does not contain any Battleneck-Slice, $|R_{0}|\leq 2^{\alpha nw}$ and
@@
First, let us note that $R_{0}$ is contained in a discrete $(2^{(1+\epsilon_{1}+\epsilon_{2})\alpha n})$-box $V_{big}$. This is actually the case because $|R_{0}|\leq 2^{\alpha nw}$ and $R_{0}$ does not contain any Battleneck-Slice, so at each coordinate it takes at most:
@@
$\left(\left\lceil\frac{2^{(1+\epsilon_{1}+\epsilon_{2})\alpha n}}{2^{\alpha n}}\right\rceil\right)^{w}=(\lceil 2^{(\epsilon_{1}+\epsilon_{2})\alpha n}\rceil)^{w}\leq(2^{(\epsilon_{1}+\epsilon_{2})\alpha n}+1)^{w}\leq(2^{(\epsilon_{1}+\epsilon_{2})\alpha n+1})^{w}$
@@ 2.3 Proof of Lemma 4
$\begin{split}\textrm{CondDeg}_{\alpha}(\pi)&\leq\log_{q}(q^{3-\delta}+2^{-\alpha\delta n}\cdot q^{3})=\\ &=\log_{q}(q^{3-\delta}+q^{-\delta}\cdot q^{3})=\log_{q}(2\cdot q^{3-\delta})=\\ &=\log_{q}{q^{3-\delta+\frac{1}{\log q}}}=3-\delta+\frac{1}{\log q}=\\ &=3-c,\end{split}$
@@
$\delta\geq c\geq\delta/2.$
@@
$\frac{1}{\alpha n}\leq\delta/2,$
@@
$n\geq\frac{2}{\delta\alpha}=O\left(\frac{1}{\epsilon^{2}}\right),$
@@ 2.4 Final Proof of Theorem 1
Now let $w\geq 3$ and the permutation $\pi^{(w)}:\{0,1\}^{wn}\times\{0,1\}^{wn}$ be defined as follows (for now let’s assume that $w$ is a multiple of $3$):
@@
It remains to estimate $\textrm{CondDeg}_{\alpha}(\pi^{(w)})$ of the presented construction. For this purpose, we consider two arbitrary $(2^{\alpha n})$-boxes: $U=U_{1}\times\ldots\times U_{w}$ and $V=V_{1}\times\ldots\times V_{w}$ (i.e. $|U_{1}|=\ldots|U_{w}|=|V_{1}|=\ldots|V_{w}|=2^{\alpha n}$) and we estimate $|\pi^{(w)}(U)\cap V|=(2^{\alpha n})^{\textrm{CondDeg}_{\alpha}(\pi^{(w)})}$.
@@
Before we move on to the main estimation, let us just recall that for any $X,Y,Z,A,B,C\subset\{0,1\}^{n}$ such that $|X|=|Y|=|Z|=|A|=|B|=|C|=2^{\alpha n}$ we have (from Lemma 4 and again the probabilistic interpretation of the inequality $\textrm{CondDeg}_{\alpha}(\pi)\leq 3-c$):
$P_{(x,y,z)\leftarrow X\times Y\times Z}(\pi(x,y,z)\in(A,B,C))\cdot(2^{\alpha n})^{3}\leq(2^{\alpha n})^{3-c}.$ (7)
@@
Finally, from (8) we have that $\textrm{CondDeg}_{\alpha}(\pi^{(w)})\leq w-\left\lfloor\frac{w}{3}\right\rfloor c$, which completes the proof.
@@ 3.1 Wider Range of Parameters
The main result (Theorem 1) largely relies on Lemma 3.14 of [6] (i.e. Lemma 3 in this paper). It contains restrictions that – we suspect – can be avoided. First of all, we ask: whether the same result can be obtained for $n$, which are not prime numbers, and also whether the parameters $n$ and $c$ must really depend on $\epsilon$ (the work of [6] even gives hope for removing these dependencies – one would like to simply use Lemma 3.1 from that paper, but unfortunately in our solution it turned out that we were able only to use Lemma 3.14, which is a version of Lemma 3.1 with weaker assumptions and a weaker thesis.).
@@ 3.3 Always-Condenser should be better then Somewhere-Condenser
Recall that our solution for the dimension $3$ is the following:
@@
While we can easily believe that the third coordinate of $\pi_{1}$ really behaves very randomly, since it is a condenser, it seems intuitive that the following construction should have even lower conductance then $\pi_{1}$ (i.e. both the second and third coordinates are condensers):
@@
However, we cannot prove formally that $\pi_{2}$ indeed is better.
@@
The above question can be naturally generalized for higher dimensions. In other words: we hypothesize that if all coordinates (and not, for example, every third one like in our construction) of a permutation are condensers, then such a permutation has an even better conductance degree than the construction that derives from Lemma 2.
@@ A.1 Theorem 3, Appendix B from [11]
$\textrm{CondDeg}_{\alpha}\leq 1+\frac{\log(3nw)}{\alpha n}\text{, provided }\alpha\leq\frac{1}{2}-\frac{1}{nw},$
@@
$\textrm{Cond}_{q}\leq 2nqw+1\text{, provided }q^{2w}\leq 2^{nw}/4.$
It is of course a consequence of the original result, because on the one hand we have (equivalence of assumptions):
$q^{2w}\leq 2^{nw}/4\iff 2^{2\alpha nw}\leq 2^{nw-2}\iff 2\alpha nw\leq nw-2\iff\alpha\leq\frac{1}{2}-\frac{1}{nw},$
@@
$\begin{split}\textrm{CondDeg}_{\alpha}&=\log_{q}(\textrm{Cond}_{q})\leq log_{q}(2nqw+1)\leq log_{q}(3nqw)=\\ &=1+\log_{q}(3nw)=1+\frac{\log(3nw)}{\log(q)}=1+\frac{\log(3nw)}{\alpha n}.\end{split}$
A.2 Main results of the paper in the notation from the paper [11]
@@
Theorem 1’.
If $n>n_{0}$ is a prime number and $\epsilon\leq\frac{\log q}{n}\leq 0.8$, then there exists a permutation $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ such that:
$\textrm{Cond}_{q}(\pi)\leq q^{w-\left\lfloor\frac{w}{3}\right\rfloor c},$

### v2

@@ 2.2 Proof of Theorem 2
We are to show that $\pi$ is a $(\alpha,\epsilon_{1},r)$-Multi-Source-Somewhere-Condenser. By definition, we should consider some independent random variables $U_{1},\ldots,U_{w}$ such that their min-entropy is at least $\alpha n$ and show that the distribution $\pi(U_{1},\ldots,U_{w})$ has the appropriate properties. However, without loss of generality (using Lemma 1), we can assume that the distribution of $U=U_{1}\times\ldots\times U_{w}$ is flat, that is, it is really just a uniformly distributed discrete $(2^{\alpha n})$-box $U$ (as defined in Definition 2). Then, after translating to the discrete interpretation, our original statement becomes:
@@
  • $|R|\xleq r\cdot 2^{\alpha nw}$, where $r=w2^{-\alpha n\epsilon_{2}}+2^{-\alpha nw\epsilon_{3}}$;
@@
    $|C_{i}\cap\{0,1\}^{n}\times\ldots\times\underset{i\text{th-position}}{\{y\}}\times\ldots\times\{0,1\}^{n}|\xleq 2^{-(1+\epsilon_{1})\alpha n}\cdot|C_{i}|.$
@@
We say that the set $X$ has at point $y\in\{0,1\}^{n}$, an $i$-Bottleneck-Slice
@@
Intuitively: $i$-Bottleneck-Slice at point $y$ means that the projection of $X$ onto the $i$-coordinate at point $y$ is nonempty, but suitably thin.
Remark: We will sometimes also say that $X$ contains a Bottleneck-Slice (without specifying the coordinate $i$). This of course means that $X$ contains an $i$-Bottleneck-Slice for some $i$.
@@ 2. Step 2:
    If $\tilde{R}$ contains any (for any $i$ and at any point $y$) $i$-Bottleneck-Slice $B$, perform the assignments:
@@ 3. Step 3:
    Loop Step 2 until $\tilde{R}$ no longer contains any Bottleneck-Slice.
@@ 6. Step 6:
Intuitively: from the set $\pi(U)$ we cut off all Bottleneck-Slices one by one until we are left with $R_{0}$, which no longer contains sets of this type. Then all cut off $i$-Bottleneck-Slices concerning the $i$-th coordinate together form $C_{i}$, unless their measure is sufficiently small, in which case $C_{i}$ is empty and these Bottleneck-Slices go to $R_{1}$. The final $R$ is the sum of $R_{0}$ and $R_{1}$.
@@
since every such set as above on the left is either empty or (as follows from the construction) equal to exactly one $i$-Bottleneck-Slice cut off from $\tilde{R}$.
@@
We have to show $|R|\xleq r\cdot 2^{\alpha nw}\text{, where }r=w2^{-\alpha n\epsilon_{2}}+2^{-\alpha nw\epsilon_{3}},$ that is,
$|R|\xleq w2^{(w-\epsilon_{2})\alpha n}+2^{(1-\epsilon_{3})\alpha nw}.$
@@
$\begin{split}|R_{1}|&\xleq w\cdot 2^{(w-\epsilon_{2})\alpha n},\\ |R_{0}|&\xleq 2^{(1-\epsilon_{3})\alpha nw}.\end{split}$
@@
The last inequality follows directly from the fact that $R_{0}$ does not contain any Bottleneck-Slice, the bound $|R_{0}|\xleq|\pi(U)|=2^{\alpha nw}$ and the property (5) (this property holds for $\pi(U)$, so obviously also for $R_{0}\subset\pi(U)$), which, however, requires a slightly deeper analysis, which (thus concluding the entire proof) we present below:
@@
If the set $R_{0}$ does not contain any Bottleneck-Slice, $|R_{0}|\xleq 2^{\alpha nw}$ and
@@
First, let us note that $R_{0}$ is contained in a discrete $(2^{(1+\epsilon_{1}+\epsilon_{2})\alpha n})$-box $V_{big}$. This is actually the case because $|R_{0}|\xleq 2^{\alpha nw}$ and $R_{0}$ does not contain any Bottleneck-Slice, so at each coordinate it takes at most:
@@
$\left(\left\lceil\frac{2^{(1+\epsilon_{1}+\epsilon_{2})\alpha n}}{2^{\alpha n}}\right\rceil\right)^{w}=(\lceil 2^{(\epsilon_{1}+\epsilon_{2})\alpha n}\rceil)^{w}\xleq(2^{(\epsilon_{1}+\epsilon_{2})\alpha n}+1)^{w}\xleq(2^{(\epsilon_{1}+\epsilon_{2})\alpha n+1})^{w}$
@@ 2.3 Proof of Lemma 4
$\begin{split}\textrm{CondDeg}_{\alpha}(\pi)&\xleq\log_{q}(q^{3-\delta}+2^{-\alpha\delta n}\cdot q^{3})=\\ &=\log_{q}(q^{3-\delta}+q^{-\delta}\cdot q^{3})=\log_{q}(2\cdot q^{3-\delta})=\\ &=\log_{q}{q^{3-\delta+\frac{1}{\log q}}}=3-\delta+\frac{1}{\log q}=\\ &=3-c,\end{split}$
@@
$\delta\xgeq c\xgeq\delta/2.$
@@
$\frac{1}{\alpha n}\xleq\delta/2,$
@@
$n\xgeq\frac{2}{\delta\alpha}=O\left(\frac{1}{\epsilon^{2}}\right),$
@@ 2.4 Final Proof of Theorem 1
Now let $w\xgeq 3$ and the permutation $\pi^{(w)}:\{0,1\}^{wn}\times\{0,1\}^{wn}$ be defined as follows (for now, assume that $w$ is a multiple of $3$):
@@
From the very construction it is obvious that $\pi^{(w)}$ is a permutation, so it remains to estimate $\textrm{CondDeg}_{\alpha}(\pi^{(w)})$ of the presented construction. For this purpose, we consider two arbitrary $(2^{\alpha n})$-boxes: $U=U_{1}\times\ldots\times U_{w}$ and $V=V_{1}\times\ldots\times V_{w}$ (i.e. $|U_{1}|=\ldots|U_{w}|=|V_{1}|=\ldots|V_{w}|=2^{\alpha n}$) and we estimate $|\pi^{(w)}(U)\cap V|=(2^{\alpha n})^{\textrm{CondDeg}_{\alpha}(\pi^{(w)})}$.
@@
Before we move on to the main estimation, let us just recall that for any $X,Y,Z,A,B,C\subset\{0,1\}^{n}$ such that $|X|=|Y|=|Z|=|A|=|B|=|C|=2^{\alpha n}$ we have (from Lemma 4 and again the probabilistic interpretation of the inequality $\textrm{CondDeg}_{\alpha}(\pi)\xleq 3-c$):
$P_{(x,y,z)\leftarrow X\times Y\times Z}(\pi(x,y,z)\in(A,B,C))\cdot(2^{\alpha n})^{3}\xleq(2^{\alpha n})^{3-c}.$ (7)
@@
Finally, from (8) we have that $\textrm{CondDeg}_{\alpha}(\pi^{(w)})\xleq w-\left\lfloor\frac{w}{3}\right\rfloor c$, which completes the proof.
@@ 3.1 Wider Range of Parameters
The main result (Theorem 1) largely relies on Lemma 3.14 of [5] (i.e. Lemma 3 in this paper), which contains certain restrictions that we suspect can be avoided. First of all, we ask whether the same result can be obtained for $n$ that are not prime numbers, and also whether the parameters $n$ and $c$ must really depend on $\epsilon$ (the work of [5] even gives hope for removing these dependencies – one would like to simply use Lemma 3.1 from that paper, but unfortunately in our solution it turned out that we were able only to use Lemma 3.14, which is a version of Lemma 3.1 with weaker assumptions and a weaker thesis.).
@@ 3.3 Always-Condenser should be better then Somewhere-Condenser
Recall that our solution for the dimension $3$ is given by:
@@
While we can easily believe that the third coordinate of $\pi_{1}$ really behaves very randomly, since it is a condenser, it seems intuitive that the following construction should have even lower conductance than $\pi_{1}$ (i.e. both the second and third coordinates are condensers):
@@
However, we cannot formally prove that $\pi_{2}$ is indeed better.
@@
The above question can naturally be generalized to higher dimensions. In other words: we hypothesize that if all coordinates (and not, for example, every third one like in our construction) of a permutation are condensers, then such a permutation has an even better conductance degree than the construction that derives from Lemma 2.
@@ A.1 Theorem 3, Appendix B from [11]
$\textrm{CondDeg}_{\alpha}\xleq 1+\frac{\log(3nw)}{\alpha n}\text{, provided }\alpha\xleq\frac{1}{2}-\frac{1}{nw},$
@@
$\textrm{Cond}_{q}\xleq 2nqw+1\text{, provided }q^{2w}\xleq 2^{nw}/4.$
This is, of course, a consequence of the original result, because, on the one hand, we have (equivalence of assumptions):
$q^{2w}\xleq 2^{nw}/4\iff 2^{2\alpha nw}\xleq 2^{nw-2}\iff 2\alpha nw\xleq nw-2\iff\alpha\xleq\frac{1}{2}-\frac{1}{nw},$
@@
$\begin{split}\textrm{CondDeg}_{\alpha}&=\log_{q}(\textrm{Cond}_{q})\xleq log_{q}(2nqw+1)\xleq log_{q}(3nqw)=\\ &=1+\log_{q}(3nw)=1+\frac{\log(3nw)}{\log(q)}=1+\frac{\log(3nw)}{\alpha n}.\end{split}$
A.2 Main Results of the Paper in the Notation from [11]
@@
Theorem 1‘.
If $n>n_{0}$ is a prime number and $\epsilon\xleq\frac{\log q}{n}\xleq 0.8$, then there exists a permutation $\pi:\{0,1\}^{wn}\to\{0,1\}^{wn}$ such that:
$\textrm{Cond}_{q}(\pi)\xleq q^{w-\left\lfloor\frac{w}{3}\right\rfloor c},$